An integer is seven more than another integer. Twice the larger integer is one less than the square of the smaller integer. Find the two integers.
Answer by mananth(16949)

let one integer be x
the other is x+7
..
twice the larger = 2(x+7)
..
1 less than square of the smaller
x^2-1
..
x^2-1=2(x+7)
x^2-1=2x+14
x^2-2x-1-14=0
x^2-2x-15=0
x^2-5x+3x-15=0
x(x-5)+3(x-5)=0
(x-5)(x+3)=0
x=5 OR x= -3
the integers are 5, 12
..
m.ananth@hotmail.ca